### v1

# MOWA: Multiple-in-One Image Warping Model

Abstract
While recent image warping approaches achieved remarkable success on existing benchmarks, they still require training separate models for each specific task and cannot generalize well to different camera models or customized manipulations. To address diverse types of warping in practice, we propose a Multiple-in-One image WArping model (named MOWA) in this work. Specifically, we mitigate the difficulty of multi-task learning by disentangling the motion estimation at both the region level and pixel level. To further enable dynamic task-aware image warping, we introduce a lightweight point-based classifier that predicts the task type, serving as prompts to modulate the feature maps for better estimation. To our knowledge, this is the first work that solves multiple practical warping tasks in one single model. Extensive experiments demonstrate that our MOWA, which is trained on six tasks for multiple-in-one image warping, outperforms state-of-the-art task-specific models across most tasks. Moreover, MOWA also exhibits promising potential to generalize into unseen scenes, as evidenced by cross-domain and zero-shot evaluations. The code will be made publicly available.
Keywords:
Image Warping Multiple-in-One Model Prompt Learning
Figure 1: MOWA is devised to address a variety of practical image warping tasks within a single framework, particularly in computational photography, where six distinct types of distortions are considered in this study. It also demonstrates an ability to generalize to novel scenarios, as evidenced in both cross-domain (unfamiliar domains) and zero-shot (unseen tasks) evaluations. The approach notably identifies and uses region-level and pixel-level fields, highlighted by red boxes, to accurately warp input images.
1 Introduction
Image warping is essential in the field of computational imaging and computer vision, serving as the foundation for numerous applications, including image rectification [7, 31], image rectangling [12, 23], camera calibration [13, 1, 17], and 3D reconstructions [10], etc. By enabling the manipulation of image data through processes such as scaling, rotation, and sheering, they allow for the seamless integration of diverse visual elements and the correction of optical imperfections. Moreover, image warping is indispensable in developing augmented reality (AR) and virtual reality (VR) applications, where it helps create immersive and realistic environments by accurately mapping textures and images onto 3D models.
Considering different inputs derived from different camera models or manipulation spaces, recent works integrate specific prior knowledge into their models to address the corresponding image warping tasks [23, 12, 18, 31, 9, 8, 11, 30, 40]. While these single-task approaches achieve significant progress, we found they suffer from two main limitations: (i) the lack of generalization and flexibility, which restricts their real-world applications since users are required to manually identify each input type and apply the appropriate single-task model. This process is time-consuming and challenging for non-professional users to judge. (ii) the substantial storage requirements for multiple task-specific models, which is impractical for some resource-limited platforms. Thus, it is crucial to develop a holistic framework capable of efficiently warping images from various camera models or manipulation spaces. Furthermore, many image warping tasks typically involve shared processes, such as motion estimation and content-aware perception. This indicates the possibility of developing a unified framework that incorporates these common image techniques.
In this work, we propose a Multiple-in-One image WArping model (named MOWA) [^1] We believe “multiple-in-one” is more rigorous than “all-in-one” or “universal” in our settings. to address various tasks in practice, as shown in Fig. 1. Specifically, we consider six representative types in the field of computational photography, namely stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks. Given the fact that learning different structures of motion is non-trivial in one model and motion representations differ significantly across various tasks, we propose to disentangle the motion estimation at both the region level and pixel level. In this hierarchical architecture, we first estimate the control points of the thin-plate spline (TPS) model [3] with increasing refined numbers, in which the feature maps are progressively warped and rectified. Such a representation excels in approximating complex motions at a region level and enables high flexibility to various motion structures. Subsequently, the warped feature maps are fed into the decoder to predict the residual pixel-level displacement, which further improves the warping results for each task, especially in the image boundaries and details.
To enable MOWA to explicitly discriminate diverse input types, a lightweight point-based classifier is devised. Adding an extra classification network based on the image features is a straightforward solution but brings high computation and storage costs. Noticing the motion structures in different warping tasks possess their specific distribution, we leverage the middle product of the image warping framework, i.e., region-level control points, to directly learn the task type. It achieves comparable performance while allowing significant parameter reduction compared to the image-based classifier since only a few 2D points are needed. Then, the soft task label predicted by this point-based classifier is used to modulate the feature maps in the decoder, dynamically boosting task-aware image warpings using a prompt learning module. Prompts are a set of learnable parameters that encapsulate essential discriminative information about different types of input, which empower a single model to efficiently traverse and harness its vast parameter space to accommodate various warping requirements.
In the experiments, we trained MOWA on six typical tasks for multiple-in-one image warping and demonstrated it outperforms state-of-the-art (SotA) task-specific models in most tasks. In addition, MOWA allows the ability to generalize to unseen scenes, as evidenced by cross-domain evaluation and zero-shot evaluation, indicating its robustness and adaptability across various scenarios. Our contributions can be summarized as follows:
  • We propose MOWA, which is the first practical multiple-in-one image warping framework. This proposed model, despite with an affordable model size, still evidently outperforms most SotA methods.
  • We propose to mitigate the difficulty of multi-task learning by decoupling the motion estimation in both the region level and pixel level.
  • A prompt learning module, guided by a lightweight point-based classifier, is designed to facilitate task-aware image warpings.
2 Related Work
Image warping is the process of manipulating an image to change its shape or alignment. This transformation is achieved by applying a spatial mapping function to the coordinates of the original image, resulting in a new image with altered geometry. In computational photography, image warping is a key technique for enhancing and manipulating images beyond traditional photography limits. This technique enables the creation of panoramic images [4, 19, 36], the correction of lens distortions [16, 2, 13], and the synthesis of novel views [39, 20, 5], etc. In the past few decades, warping techniques have significantly contributed to the development of advanced imaging applications beyond those mentioned above, offering greater flexibility and creativity. For example, the image boundaries can be twisted by different manipulations, leading to visually unpleasant layouts and negative effects on downstream vision tasks. Nevertheless, in practical scenarios, most users favor rectangular boundaries due to their compatibility with standard display formats, facilitating ease of sharing, printing, and publication [12, 11]. Therefore, researchers have developed diverse image rectangling methods to warp the image boundaries to be straight [12, 23, 18, 14, 37, 38]. Most of them follow the principle of content-aware image warping to avoid the large distortion on the original distribution when rectangling the image. Besides, different motion representations are also exploited, such as the mesh [12, 23] and control points [18], to formulate the warping process.
Excluding the customized manipulations, some special camera models can introduce geometric distortion onto the captured images, $e.g.$, radial distortion, rolling shutter distortion, and perspective distortion. The images’ semantic features significantly disobey the real-world rules due to those distortions. To address this issue, there is an exploration of distortion correction approaches [16, 2, 34, 7, 33, 21, 27, 30, 29, 40] aimed at warping the distorted input to a geometrically reasonable one. Particularly, regression-based methods [2, 27] learn the camera and distortion parameters from the input image and correct the distortion by simulating the imaging process of a predefined camera model. In contrast, reconstruction-based methods [16, 34, 7, 33, 30] directly learn the pixel-wise displacement between the distorted image and its ground truth, facilitating the model-free correction and enabling the end-to-end training.
The above works achieve remarkable progress on various tasks, of which well-designed network architectures and tailored motion representations are studied. However, they need to train an individual model for each specific warping type and require the prior knowledge of the camera model or customized manipulations. In this work, we propose a multiple-in-one framework to involve these typical and practical image warping tasks.
3 Multiple-in-One Warping Model
3.1 Problem Definition
In this study, we consider six representative and practical image types in the field of computational photography, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos. These types are further classified into two groups. The first four types (stitched images, rectified wide-angle images, unrolling shutter images, and rotated images) struggle with irregular boundaries as the original images are manipulated by some customized operations, such as image stitching, distortion correction, and rotation. Therefore, image rectangling is proposed to reshape these irregular boundaries while keeping the distribution of content unchanged. The last two types (fisheye images and portrait photos) show inherently geometric distortions imaged by special camera models, such as radial distortion in fisheye images and perspective distortion in portrait photos. Correcting these distortions is crucial to scene understanding and aesthetic appreciation.
Figure 2: Overview of the proposed multiple-in-one image warping model (MOWA).
Figure 2 shows the overall framework of the proposed MOWA. It takes the image and mask as input and estimates the TPS control points with increasingly refined numbers. In this region-level motion estimation, the feature maps are progressively warped and rectified. Subsequently, the warped features are fed into the decoder to predict a residual pixel-level motion. To enable task-aware and expandable capabilities, a lightweight point-based classifier and prompt learning module are designed. We elaborate on the details of each module as follows.
3.2 Motion Estimation Module
Learning multiple warping types in one model is challenging since the network needs to balance different complexities of the multiple motion types in motion estimation. Furthermore, the model’s scalability would be restricted if the motion representation is hand-crafted for specific tasks. Hence, we propose a flexible and hierarchical architecture for general image warping in MOWA. As shown in Fig. 2, the motion estimation is disentangled at both the region level, where the number of TPS control points progressively increases, and the pixel level, where a residual map is predicted to further compensate the estimated TPS flow.
The TPS transformation [3] stands out for its remarkable ability to model complex motions [6, 28, 15]. It is adept at performing image warping based on two sets of region-level control points, namely $\bm{Q}=[\bm{q}_{1},\bm{q}_{2},\cdots,\bm{q}_{N}]^{T}\in\mathbb{R}^{N\times 2}$ for the source image and $\bm{Q}^{{}^{\prime}}=[\bm{q}^{\prime}_{1},\bm{q}^{\prime}_{2},\cdots,\bm{q}^{\prime}_{N}]\in\mathbb{R}^{N\times 2}$ for the target image. To minimize the distortion of the source and target images, an energy term is introduced to penalize the Euclidean distance between the transformed source points $\mathcal{T}(\bm{q}_{i})$ and the target points $\bm{q}^{\prime}_{i}$, i.e., $\sum_{i=1}^{N}\lVert\mathcal{T}({\bm{q}_{i}})-{\bm{q}^{\prime}_{i}}\rVert_{2}^{2}$. This penalty results in a spatial deformation function parameterized by the control points, effectively capturing the intricate deformations across the image and maintaining the overall structural integrity. Specifically, the derived spatial deformation function can be expressed as follows:
$\mathcal{T}(\bm{q})=\bm{A}\begin{bmatrix}\bm{q}\\ 1\end{bmatrix}+\sum_{i=1}^{N}U\left({\left\lVert\bm{q}^{\prime}_{i}-\bm{q}\right\rVert}_{2}\right)\bm{w}_{i},$ (1)
where $\bm{q}$ represents a point located in the source image. $\bm{A}\in\mathbb{R}^{2\times 3}$ and $\bm{w}_{i}\in\mathbb{R}^{2}$ are the transformation parameters, $U(\cdot)$ is a radial basis function to quantify the influence of the control point, more details can be found in literature [3]. Notably, this deformation function plays a key role in determining the deformation induced by each control point, thereby shaping the overall transformation.
Motion estimation acts as the fundamental stage in image warping, presenting particular challenges in the context of multi-in-one task learning. To enhance the capability of our model in motion estimation, we design a progressive motion estimation module. More specifically, this module cascades a sequence of TPS transformation heads that gradually increase the number of control points. The control points predicted by the preceding head are upsampled and integrated into the prediction of the next head. Subsequently, these control points are arranged to generate a mesh. Then we adopt the TPS transformation to warp this mesh, aiming to align it with the regular mesh defined on the ground truth image. In the implementation, considering the cascade of fully connected layers introduces significant computation and storage costs, we use one or two convolution layers to predict the control points after each TPS transformation head. The pipeline of cascaded TPS transformation heads can be expressed by:
$\bm{q}^{(t)}=h^{(t)}[\mathcal{R}(\mathcal{F}^{(t-1)},\bm{q}^{(t-1)})]+\text{UP}[\bm{q}^{(t-1)}],$ (2)
where $h^{(t)}$ is the $t$-th TPS transformation head, $\mathcal{F}^{(t-1)}$ and $\bm{q}^{(t-1)}$ are the feature map and control points of the $(t-1)$-th head, respectively. $\mathcal{R}(\cdot,\cdot)$ represents the warping operation for feature maps given control points, and $\text{UP}[\cdot]$ is a customized upsampling layer for control points.
While TPS transformation is flexible to various tasks, it suffers from a limited degree of freedom to describe the detailed motions. To alleviate this limitation, we further complement the region-level motion representation with a pixel-level residual flow. Specifically, we first rectify the feature map $\mathcal{F}^{(T)}$ using the corresponding control points $\bm{q}^{(T)}$ in the last transformation head, and then feed the rectified feature map into a decoder network $f(\cdot)$ to predict the desirable residual flow. Like common U-Net architectures, the shallow features in the encoder are transited into the decoder using skip-connection. To eliminate the blur effect by multiple warpings (interpolation operation involved), we densify the TPS control points to pixel level and couple it with the residual flow to directly warp the input image. The final warping result can be formulated as follows:
$I^{\prime}=W(f[\mathcal{R}(\mathcal{F}^{(T)},\bm{q}^{(T)})]+\text{DE}[\bm{q}^{(T)}],I),$ (3)
where $W(\cdot,\cdot)$ denotes the warping operation given the flow map and input image, $\text{DE}[\cdot]$ densifies the sparse control points to a dense flow map, which can be regarded as a special case of TPS upsampling layer $\text{UP}[\cdot]$.
Figure 3: Motion structures in different tasks possess their specific distribution, which potentially exists in a 2D point space. Discriminating these motion structures as a classification task can also help the image warping performance as exhibited in visual comparisons.
3.3 Point-based Task Classifier
It is straightforward to design a task classifier in terms of the input image. However, such a design brings an unavoidable issue of high computation complexity due to the redundant image features. Instead, we propose a lightweight task classifier based on the TPS points predicted by the motion estimation module. Our motivation stems from the fact that the motion structures in different tasks possess their specific distribution, which potentially exists in a 2D point space as shown in Fig. 3. To this end, we design a PointNet-like network [26] to predict the task type. Specifically, it takes the local coordinates of motion with the global image features (after maxpooling) by point-wise concatenation along the last dimension as input and outputs the soft task label, as shown in Fig. 2.
In addition to the task classification function, our point-based task classifier can further improve image warping performance. This improvement is due to the high-level guidance provided by the task classifier to the motion estimation module through gradient back-propagation. Figure 3 (right) depicts a typical example, in which the fisheye rectification result shows a less distorted shape, and the predicted control points of the stitched image are more tightly aligned to the image boundary. Experiments demonstrate our point-based task classifier achieves comparable results while having less than ×$\times 50$ parameters compared with the image-based classifier. More quantitative results are presented in Sec. 4.3.
3.4 Prompt Learning Module
Once the inputs are classified by the proposed point-based network, we leverage the predicted soft task label to modulate the feature maps in our MOWA. In particular, a prompt learning block is inserted into each layer in the decoder as a plug-and-play module. Prompt learning aims to tackle the challenge of generalizing in various image warping tasks by aiding the network in comprehending the specific task at hand. The prompts serve as a flexible and lightweight component to encode motion context across multiple scales within the image warping network.
Assuming the task number is $N$, we introduce a set of learnable parameters as our prompts, namely $\{\bm{P}_{i}\}_{i=1}^{N}$. By denoting the predicted soft label by the task classifier as $\bm{\Phi}$, we modulate the feature maps $\mathcal{F}$ in the decoder network by the prompts as follows:
$\mathcal{F}^{m}=\text{Conv}_{1\times 1}\big{(}\mathcal{F}\oplus\sum_{i=1}^{N}{\Phi_{i}\bm{P}_{i}}\big{)},$ (4)
where $\oplus$ represents the concat operation, $\text{Conv}_{1\times 1}$ is a convolution layer with $1\times 1$ kernel size aiming to reduce the channel dimension of concatenated features. By integrating prompts with the feature of MOWA, we can effectively enrich the representations with task-specific knowledge. This adaptive process empowers the network to efficiently warp images by dynamically influencing its behavior.
3.5 Training Loss
After predicting the TPS control points and the residual flow, the warped image can be obtained by Eq. (3). Following previous works [23, 18], we first exploit three losses to train our multiple-in-one image warping framework, e.g., image reconstruction loss $\mathcal{L}_{Rec}$, perceptual loss $\mathcal{L}_{Per}$, and inter-grid loss $\mathcal{L}_{Grid}$. The reconstruction loss and perceptual loss supervise the warped image at pixel level and feature level, respectively, while the inter-grid loss constrains the edges of two consecutive deformed grids to be co-linear. Considering the ground truth of warping flow is available in the dataset of portrait photos, we also add the reconstruction loss $\mathcal{L}_{Flow}$ on the predicted flow of the portrait correction task. Moreover, we provide middle-level supervision on the warped results from the TPS prediction heads with a set of exponentially growing weights. To train the point-based task classifier, the standard cross-entropy loss $\mathcal{L}_{Cls}$ is applied. Overall, the final loss can be expressed by:
$\mathcal{L}=\underbrace{\mathcal{L}_{Rec}+\mathcal{L}_{Per}+\mathcal{L}_{Grid}+\lambda_{Flow}\mathcal{L}_{Flow}}_{Image\ Warping}+\underbrace{\lambda_{Cls}\mathcal{L}_{Cls}}_{Task\ Classifier},$ (5)
where $\lambda_{Flow}$ and $\lambda_{Cls}$ are the hyper-parameters to balance different losses, both of them are empirically set to $0.1$.
4 Experiments
To demonstrate the effectiveness of the proposed multiple-in-one image warping method, we evaluate its performance on six representative distorted types, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos.
4.1 Experimental Settings
Implementation Details. We train the proposed model using the Adam optimizer with the momentum terms of $(0.9,0.999)$ on 8 NVIDIA A100 GPUs. The learning rate starts with a linear warm-up in the first three epochs and then decays from $1e^{-4}$ to $1e^{-6}$ following a cosine schedule in the remaining epochs. The batch size is set as 64. For the network architecture, the Transformer blocks with shifted windows [22, 32] are used in both the encoder and decoder except for the input projection layer and output projection layer. In TPS prediction heads, we adopt the convolution layers with different kernels to predict increasing numbers of control points, and the numbers are set to $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$. More details of our network can be found in supplementary material. The complete framework is trained with a fixed input size of $256\times 256$. At the first 10 epochs, we solely train and supervise the TPS prediction heads with the point-based task classifier. Afterwards, all modules are trained collectively. During inference, MOWA supports the image warping for any resolution by scaling the predicted TPS control points and residual flow.
Datasets. We use the public benchmarks from recent SotA works, including the image rectangling datasets [23, 24, 18] and the distortion correction datasets [34, 30, 40]. Since there is no available training dataset for unrolling shutter image rectangling, we use the rolling shutter correction dataset [33] and follow the standard data construction process from previous methods [23] to synthesize the paired data. This dataset will also be made public.
Metrics. Following previous works, we select PSNR and SSIM as metrics to quantitatively measure the quality of the warped results. In particular, for the portrait photos, the ShapeAcc metric is applied as suggested in Tan et al. [30]. It is specially designed for the quality of face correction, which calculates the similarity between corrected portraits and the stereographic projection of its original input.
4.2 Comparison Results
We compare the proposed MOWA with recent SotA methods on each task, including Deep_Rect [23], He et al. [12], RecRec [18], Feng et al. [7], Shih et al. [29], Tan et al. [30], and Zhu et al. [40].
Figure 4: Qualitative comparison of our multiple-in-one framework MOWA to the SotA image warping models. The red dotted lines mark the horizon. The arrows highlight the inferior warped parts such as the irregular boundaries and distorted semantics.
Qualitative Comparison. As shown in Fig. 4, we visualize the comparison results of different methods on the testing datasets. These qualitative results demonstrate that our multiple-in-one method can handle well across various tasks, scenes, and resolutions, compared with the SotA methods specially designed for each single task. For example, for the rotated images, our method can rearrange the input to a rectangle one while keeping the original geometric layout reasonable. On the contrary, distorted buildings can be observed in the results of previous works [12, 23], in which the physical world rules such as the horizon are also perturbed. For other rectangling tasks like the stitched image, unrolling shutter image, and rectified wide-angle image, our method shows a better visual appearance especially in the image boundaries, allowing promising structural integrity among the comparison methods. In some challenging cases such as the first and second rows in stitched images, the image boundaries are dramatically stretched, but our method can still warp the images to expected structures. One important reason is that MOWA learns the generalized warping strategy from different tasks since it can extract some common knowledge from them. In addition, our MOWA mitigates the difficulty of motion estimation by disentangling it at both the region level and pixel level. Consequently, diverse structures of motions can be progressively approximated and the image details can be preserved. For the fisheye image and portrait photos, MOWA is capable of recovering the realistic geometric distribution from the inputs, despite the radial distortion or perspective distortion.
Table 1: Quantitative evaluation of the proposed multiple-in-one framework MOWA to the SotA image warping models.
Warping Tasks Metrics
Input Type Methods PSNR $\uparrow$ SSIM $\uparrow$ ShapeAcc $\uparrow$ Parameter
Stitched Image Deep_Rect [23] 21.28 0.7140 - 52.14M
Ours 20.42 0.6307 - 49.93M
Rectified Wide-Angle Image RecRecNet [18] 18.68 0.5450 - 62.70M
Ours 18.78 0.5479 - 49.93M
Unrolling Shutter Image RecRecNet [18] 21.48 0.7602 - 62.70M
Ours 21.83 0.7798 - 49.93M
Rotated Image He et al. [12] 17.63 0.4880 - -
Deep_Rect [23] 19.89 0.5500 - 52.14M
Ours 20.83 0.5874 - 49.93M
Fisheye Image Feng et al. [7] 21.72 0.7167 - 11.65M
Ours 23.92 0.8076 - 49.93M
Portrait Photos Shih et al. [29] - - 97.253 -
Tan et al. [30] - - 97.490 -
Zhu et al. [40] - - 97.491 8.79M
Ours - - 97.475 49.93M
Quantitative Comparison. We report the quantitative evaluation results in Table 1. The proposed multiple-in-one image warping jointly learns six tasks and achieves promising performance compared with the single-task methods. For example, MOWA outperforms the SotA methods in rectified wide-angle images, unrolling shutter images, rotated images, and fisheye images. Although the purposes and motion structures of these tasks differ from each other, the proposed method can enable a unified image warping framework with an elaborately designed hierarchical motion estimation architecture and task-aware prompt learning strategy. Moreover, MOWA achieves comparable image warping performance for the stitched image and portrait photos, without intolerable performance degradation when involving more tasks and data. However, previous methods [18, 12, 7, 23, 30] tailor the specific knowledge into their models to address the single image warping task, lacking the generalization and flexibility to real-world applications.
Computation Complexity Comparison. In Table 1, we also compare the computation complexity of previous methods which have available models, and our MOWA. The results demonstrate our model size is reasonable as a multiple-in-one image warping framework. Even compared to the SotA models designed for the specific task [23, 18], our MOWA has fewer parameters to achieve better or comparable warping performance. The underlying reason is that the proposed motion estimation module discards the heavy fully connected layers and replaces them with convolutional layers. Then, the predicted region-level TPS points are further compensated with the pixel-level displacement from a compact convolutional decoder. Besides, the shared knowledge across different tasks can also help to decrease the parameter requirements of a multi-task model.
Figure 5: Ablation study on the proposed motion estimation module. The predicted TPS control points are shown with the size of $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$, from the left to right. The coarse results and final results are obtained by warping the input using the first control points and final flow (coupled with the last TPS points and residual flow), respectively.
4.3 Ablation Study
Considering the aim of a multiple-in-one framework is to achieve holistic performance across various tasks, we mainly compare the different variants of the framework in terms of the average warping metrics. Additionally, the same image quality metrics (PSNR and SSIM) are shared in the first five tasks, but the portrait correction task has its own metrics like ShapeAcc. Thus, the average PSNR and SSIM from the first five tasks are reported in this part.
Table 2: Ablation study on the proposed motion estimation module. The baseline represents the predicted control points with a size of $12\times 12$. “10x10x10x10” means 4 heads are applied and each head predicts the control points with a size of $10\times 10$. Other settings are also presented in this form. “Ours” denotes the combination of “10x12x14x16” and residual flow.
Metrics Baseline 12x12x12x12 14x14x14x14 16x16x16x16 10x12x14x16 Ours
PSNR 20.29 20.38 20.42 20.02 20.48 20.84
SSIM 0.6279 0.6311 0.6406 0.6148 0.6418 0.6572
Motion Estimation. It is challenging to estimate multiple motions in one model since the motion’s complexities and patterns significantly differ across various tasks. For this purpose, we propose a flexible and hierarchical architecture to disentangle the motion estimation at the region level and pixel level. As shown in Fig. 5, better localization performance of the formed mesh can be achieved by increasing the number of TPS points. Besides, the pixel-level residual flow can provide a higher degree of freedom for the motion than only the region-level motion representation, improving the warping results, particularly in the image boundaries and details. Table 2 quantitatively demonstrates the effectiveness of the proposed hierarchical motion estimation module. We also found the upper bound occurs when increasing the number of control points, which suggests the performance of multiple-in-one image warping would be limited without proper decoupling motion manner.
Table 3: Ablation study on different task classifiers for the multiple-in-one image warping framework.
Metrics w/o Classifier Image-based Classifier Point-based Classifier Ours
PSNR 20.48 20.58 20.63 20.83
SSIM 0.6418 0.6451 0.6463 0.6558
Parameters - 3.39M 0.0592M 0.0594M
Task Classifier. As the image-based classifier involving redundant image features burdens the main image warping framework, we propose a lightweight point-based classifier to learn the task type from each input image. As listed in Table 3, four baseline models are designed: image warping model without the task classifier, with the image-based classifier, with the point-based classifier, and with the point-based classifier compensated with the pooling global features (Ours). Note that we validate different task classifiers only with the TPS prediction modules, showing their direct influence on the region-level motion estimation. The quantitative results demonstrate that we can obtain evident performance gain beyond the vanilla network by adding the task classifier, indicating that the information on task type is meaningful in multi-task training.
Prompt Learning. In our MOWA, a prompt learning module is designed to modulate the feature map with the soft task label, which helps to dynamically navigate its extensive parameter space to achieve task-aware image warping. By combining this module into MOWA, the averaged PSNR metric of all warping results gains $+0.21$dB improvements beyond the baseline. More experimental results can be found in supplementary materials.
Figure 6: Generalization evaluation of the proposed MOWA: cross-domain evaluation (top) and zero-shot evaluation (bottom). The visualized flow and control points are both predicted by MOWA. In image retargeting results, the red dotted lines measure the stretch extent of face or body by warping operations.
4.4 Generalization
We show the generalization ability of the proposed MOWA in terms of the cross-domain and zero-shot evaluations in Fig. 6. For the cross-domain evaluation, the new inputs belong to the above six practical image warping tasks but they are captured in real-world settings with different cameras, resolutions (can be up to 4K), and scenes. For the zero-shot evaluation, we consider a new image warping task, i.e., image retargeting, which aims to flexibly change the image scale without distorting the content as much as possible.
The visualization results demonstrate that MOWA can well extend to real-world scenarios though its training datasets are mostly synthesized by hand-crafted camera models or manipulation spaces. One possible reason is the multiple-in-one model can naturally address the overfitting issue on specific datasets by learning various tasks. In addition to the cross-domain evaluations, we find that while our model does not involve the image retargeting task during training, it is still able to warp the image based on the “content-aware” principle. As we can observe, the predicted control points are accurately aligned to the image boundaries. Such knowledge transferring to new tasks potentially benefits from the shared motion perception across different tasks. Therefore, our results show fewer geometrical distortions for the foreground than the crop and resize operation, with the face and body experiencing less stretching.
5 Conclusion
We have proposed MOWA in this work, the first multiple-in-one image warping framework in the field of computational photography. It considers six representative and practical tasks in one learning model and uses a unified motion representation to achieve various warping purposes. In particular, to mitigate the difficulty of approximating diverse motions of different tasks, we propose to disentangle the motion estimation at both the region level and pixel level. Then, we enable MOWA the explicit task-aware ability by introducing a lightweight point-based classifier. Compared to the common image-based classifier, it can achieve comparable performance while offering significant parameter reduction. Subsequently, we feed the soft task label predicted by the task classifier into a prompt learning module and further modulate the feature maps in the decoder, which facilitates a single model to efficiently navigate and leverage its extensive parameter space to meet various warping requirements. Comprehensive experiments demonstrate that MOWA outperforms different SotA methods specifically designed for each single task, with an affordable model size.
In the future, we plan to empower MOWA with cross-view and cross-modal abilities, aiming to build a foundation model for universal image warping.
Appendix 0.A Supplementary Materials
0.A.1 Overview
In this document, we provide the following supplementary content:
  • More implementation details of the multiple-in-one framework (Section 0.A.2).
  • More results and analysis of ablation study (Section 0.A.3 and Section 0.A.4).
  • More qualitative results of the comparison methods (Section 0.A.5).
  • More generalization evaluations (Section 0.A.6).
  • Limitation discussion (Section 0.A.7).
0.A.2 More Implementation Details
We design the architecture of our MOWA based on the encoder-decoder style, it enables both region-level control point regression and pixel-level residual flow prediction. To be more specific, the Transformer blocks with shifted windows (with the size of 8) are exploited in each layer except for the input and output projection layers. The basic dimension of channels is set to 32 and linearly increases along the layers in the encoder network, which is oppositely decreased to 2 in the decoder network. During training, all inputs are sized to $256\times 256$ and the size of the last feature maps in the encoder network is $16\times 16$. Moreover, the depths of each Transformer block are set to 2 and the head numbers of multi-head self-attention are $[1,2,4,8,16,16,8,4,2]$ along the whole layers. To establish a compact network, the convolutional layers are used to regress the coordinates of 2D control points, which enable significant parameter reduction compared with the fully connected layers. The configuration details of these regression heads are listed in Table 4. For the lightweight point-based classifier, three 1D convolutional layers with channel dimensions of $256,256,512$ are used to extract the features of input points and then three fully connected layers with unit numbers of $512,256,6$ are used to classify their task types.
In this work, our primary objective is to contribute the first multiple-in-one image warping model to the computational photography community. The aforementioned configuration has been proven effective in simultaneously learning six tasks within a single model. It achieves superior or comparable performance to the state-of-the-art single-task methods even with fewer parameters. We believe more complex architectures could further improve the image warping performance and leave it as one of the future works.
Table 4: The configurations of convolutional layers to regress different sizes of control points (the size of input feature maps is $16\times 16$).
Configuration $8\times 8$ $10\times 10$ $12\times 12$ $14\times 14$ $16\times 16$
Kernel Size $3\times 3$$5\times 5$, $3\times 3$} $5\times 5$ $3\times 3$ $3\times 3$
Stride 2 {1, 1} 1 1 1
Padding 1 {0, 0} 0 0 1
0.A.3 More Analysis and Ablations on Multi-Task Learning
Figure 7: Evaluations of the multi-task learning and effectiveness of the proposed prompt learning module. The normalized PSNR and SSIM of validation data are visualized for six tasks. We recap different task types at the bottom.
To intuitively analyze the influence of multi-task learning and the effectiveness of the proposed prompt learning, we visualize the quantitative metrics of the warping results of validation images for each task. As illustrated in Figure 7, the values of normalized PSNR and SSIM are plotted along different training epochs. Considering the ranges of PSNR and SSIM of different warping tasks significantly differ from each other, we normalize all values into the range of $[0,1]$ to eliminate the data bias. For the portrait photos, we leverage the available warping flow in its training dataset and obtain the corrected images to compute the PSNR and SSIM. For other tasks, we directly use the ground truth of warped images in the corresponding datasets.
From Figure 7 (a), we have the following three observations: (1) Different tasks show various levels of difficulty when training a multiple-in-one image warping model. For example, learning to warp the unrolling shutter image (task 3) is generally easier than other tasks, which shows the fastest convergence at the first 80 epochs. The reason is that the structures of unrolling shutter images are basically regular, where more than two boundaries are straight and do not need to be rectified. On the contrary, learning to warp the stitched images (task 1), rectified wide-angle images (task 2), and portrait photos (task 6) is more challenging (slow convergences can be observed) since their boundaries or warping motions vary greatly in datasets. Especially in the portrait photo dataset, the numbers, shapes, and locations of the faces are quite diverse. (2) The multiple-in-one model tends to sacrifice the performance of some individual tasks to achieve the improvement of holistic warping performance. Particularly, the accuracy of warping the unrolling shutter image is dramatically reduced to improve the warping performance of other tasks since the $80$th epoch. Such a trade-off among different tasks facilitates an overall improvement but leads to performance conflicts between some tasks. (3) The relationship of different tasks can be positive or negative. For example, the curves of the stitched image and rotated image show consistent trends during the MOWA’s training as they share the similar warping principle, i.e., rectangling the irregular image boundaries and keeping the content unchanged. Thus, meaningful interactions between these two tasks could happen if learning a multiple-in-one model. However, the curve of warping the fisheye image shows a converse trend to those of the stitched image and rotated image. The reason for this is that correcting the radial distortion present in fisheye images significantly alters the scene’s layout while leaving the image boundaries almost untouched. This approach contrasts with the foundational principles of tasks based on image rectangling. As a result, the imbalanced convergences can be noticed across these tasks with negative relationships.
Figure 8: t-SNE visualization of the learned prompts in MOWA.
To dynamically boost the task-aware image warping in a single model, we propose a prompt learning module and a point-based task classifier. As shown in Figure 7 (b), the performance conflicts of different tasks are relieved by prompt learning. All tasks show a similar improvement trend as the training epoch increases, without dramatic performance degradation in certain individual tasks. More importantly, the multiple-in-one framework achieves the unified and best warping performance on all tasks at the end of training epochs. This phenomenon suggests that the framework knows to discriminate and warp different input types using the learned task-specific prior knowledge. Our designed prompts enable MOWA to efficiently traverse and harness its vast parameter space to meet various warping requirements. We also visualize the t-SNE of the learned prompts of MOWA in Figure 8. We can observe these prompts are well-clustered based on the task types, demonstrating the ability of prompts to learn discriminative motion context that aids in holistic image warping.
Figure 9: Ablation study on TPS regression heads. We show different warping performances from two groups of control points. For each comparison, we show the warping results from the control points with the size of $16\times 16$, $16\times 16$, $16\times 16$,16×$\times 16$ 16 (top), and $10\times 10$, $12\times 12$, $14\times 14$, $16\times 16$ (bottom). The arrows highlight the under-rectified distortion parts.
0.A.4 More Analysis and Ablations on Motion Estimation
To eliminate the difficulty of estimating various warping motions, we propose to disentangle the motion estimation at both the region level and pixel level. In the region-level estimation, the hierarchical TPS control points with increasing numbers are designed to gradually approximate the motion structure. As shown in Figure 9, the results of the ablation study demonstrated the proposed TPS regression heads achieve more accurate warping performance on severe geometric distortion correction, compared with the settings with even more control points. This indicates the upper bound of the control point numbers exists in certain warping tasks, where adding more control points would restrict the model’s flexibility to estimate complex motions. The MOWA, by contrast, learns the progressive warping rules in a spatially coarse-to-fine manner, allowing more flexible structures at the beginning levels of cascade regression heads.
0.A.5 More Comparison Results
We show more qualitative comparison results in Figure 10, Figure 11, Figure 12, and Figure 13. The results demonstrate the proposed MOWA shows superior performance across diverse tasks, scenes, and geometric distortions, verifying its promising practicality beyond previous methods customized to a single task.
Figure 10: More comparison results on the stitched image test dataset.
Figure 11: More comparison results on the rotated image test dataset.
Figure 12: More comparison results on the rectified wide-angle image, unrolling shutter image, and fisheye image test datasets. For each comparison, we show the input image, results of comparison methods [18, 7], MOWA, and ground truth from left to right.
Figure 13: More comparison results on the portrait photo test dataset.
0.A.6 More Generalization Evaluations
In addition to the performance on the test datasets, we demonstrate the proposed MOWA has favorable generalization ability to unfamiliar domains and unseen tasks. In particular, we show cross-domain evaluations of MOWA in Figure 14, Figure 15, Figure 16, and Figure 17. The cross-domain datasets are derived from real-world datasets [25, 18, 35, 40]. Compared to the training datasets, the images in real-world datasets are captured by various cameras or obtained by practical manipulations. As shown in these figures, while the significant domain gap exists, the proposed approach can well extend to other domains. It is noticed that MOWA also exhibits satisfied robustness to the noisy data. For instance, in Figure 17, although MOWA is trained on the synthetic dataset with clean masks, the warping results of the real-world dataset with noisy pseudo masks (the clean masks are not available in some practical applications) are still structurally reasonable.
The zero-shot evaluation is shown in Figure 19. We consider a new image retargeting task as an unseen domain to validate the MOWA’s generalization. Compared to the resize operation, MOWA can retarget the resolution of the input image with less stretching of persons and objects. Compared to the classical image rectangling method [12], fewer geometric distortions are introduced by warping the input image with MOWA.
Figure 14: Cross-domain evaluation on the real-world high-resolution stitched images. The resolution of the input images ranges from $1191\times 722$ to $3616\times 2048$.
Figure 15: Cross-domain evaluation on the real-world fisheye images.
Figure 16: Cross-domain evaluation on the real-world portrait photos.
Figure 17: While MOWA is trained on the synthetic datasets with clean masks, it shows good robustness to the real-world data with noisy pseudo masks. The motion estimation of real-world images, e.g., the predicted control points, and warping flow, exhibit structurally reasonable distributions similar to those of the synthetic images.
Figure 18: Failure cases of the proposed approach. It fails to accurately warp the input image with challenging image boundaries using a certain number of control points.
Figure 19: Zero-shot evaluation on the new task: image retargeting.
0.A.7 Limitation Discussion
We show some failure cases of MOWA in Figure 18. In these cases, we can find that image boundaries are much more complicated and rarer than most samples. Consequently, it is challenging to approximate the accurate motion structure with a certain number of control points. This limitation could be addressed by adding more control points and cascading more TPS regression heads. Besides, scaling up the resolution of the input image could potentially improve the warping performance on image boundaries and details.